1. Consider rolling two dice once. Calculate the probability of the...
A, rolling a sum of 5 or an even sum
B, rolling a sum of 7 or 6
C,rolling a sum of 10 or an odd sum

2.six(6) women, 4 men, 3 boys and 7 girls enter a competition. They each have an equally likely chance of winning. What is the probability that the winner will be..
A, female?
B, a child?
C, a man?
1
Expert's answer
2021-07-11T09:07:05-0400
  1. Rolling two dice once and summing the outcomes is summarized in the following table.


(a). "P(5\cup even)"

The two events (sum of 5 and even sum) are disjoint. Thus,

"P(5\cup even)=P(5)+P(even)"

"P(5)=\frac{4}{36}=\frac{1}{9}"

"P(even)=\frac{18}{36}=\frac{1}{2}"

"P(5\cup even)=\frac{1}{9}+\frac{1}{2}"

"=\frac{11}{18}"


(b). "P(7\cup 6)"

"P(7\cup 6)=P(7)+P(6)"

"=\frac{5}{36}+\frac{6}{36}"

"=\frac{11}{36}"


(c). "P(10\cup odd)" (disjoint)

"P(10\cup odd)=P(10)+P(odd)"

"=\frac{3}{36}+\frac{18}{36}"

"=\frac{7}{12}"


2.

a. P(female)

6 women and 7 girls out of 20

"P(female)=\frac{13}{20}"


b. P(Child)

3 boys and 7 girls out of 20

"P(Child)=\frac{10}{20}"

"=\frac{1}{2}"


c. P(man)

4 men out of 20

"P(Man)=\frac{4}{20}"

"=\frac{1}{5}"